OpenStudy (anonymous):

At a school carnival, the diameter of the mat of a trampoline is 10 feet and the diameter of its metal frame is 12 feet. What is the length, in feet, of the metal frame that surrounds the trampoline? Use 3.14 for π and round your answer to the nearest tenth.

OpenStudy (anonymous):

does anyone understand

OpenStudy (michele_laino):

I think that the requested length is given by the subsequent formula: \[\Large L = 2 \times 3.14 \times \frac{d}{2}\] where d= 12 feet

OpenStudy (anonymous):

oh

OpenStudy (michele_laino):

hint: \[\Large L = 2 \times 3.14 \times \frac{d}{2} = 2 \times 3.14 \times \frac{{12}}{2} = ...?\]

OpenStudy (rushwr):

I agree @Michele_Laino @bubbleslove1234 As they are talking about a circle kind of like thingy, when they ask the surrounding it is gonna be the circumference of the circle. So we are able to use this equation \[C= 2\pi r\]

OpenStudy (rushwr):

where as r is the radius of the circle. ! Hence we can substitute it to \[\frac{ d }{ 2}\]

OpenStudy (rushwr):

Hope u understood me